### v1

# Trajectory Balance with Asynchrony: Decoupling Exploration and Learning for Fast, Scalable LLM Post-Training

Abstract
Reinforcement learning (RL) is a critical component of large language model (LLM) post-training. However, existing on-policy algorithms used for post-training are inherently incompatible with the use of experience replay buffers, which can be populated scalably by distributed off-policy actors to enhance exploration as compute increases. We propose efficiently obtaining this benefit of replay buffers via Trajectory Balance with Asynchrony (TBA), a massively scalable LLM RL system. In contrast to existing approaches, TBA uses a larger fraction of compute on search, constantly generating off-policy data for a central replay buffer. A training node simultaneously samples data from this buffer based on reward or recency to update the policy using Trajectory Balance (TB), a diversity-seeking RL objective introduced for GFlowNets. TBA offers three key advantages: (1) decoupled training and search, speeding up training wall-clock time by 4x or more; (2) improved diversity through large-scale off-policy sampling; and (3) scalable search for sparse reward settings. On mathematical reasoning, preference-tuning, and automated red-teaming (diverse and representative post-training tasks), TBA produces speed and performance improvements over strong baselines.
Figure 1: TBA performs rapid, scalable exploration of model responses, improving RL efficiency on the GSM8K mathematical reasoning task. All plotted points use 4xA100 GPUs (or comparable L40S GPUs). DPO and RLOO baselines taken from Noukhovitch et al. (2025), PPO and VinePPO baselines taken from Kazemnejad et al. (2024). The baseline model is the SFTed RhoMath-1B (Lin et al., 2024) model, which obtains 40.3% accuracy after SFT and before RL. Appendix A has details.
1 Introduction
Post-training through reinforcement learning (RL) is a critical step in enhancing large language models (LLMs), aligning them with human preferences and improving their reasoning abilities (Christiano et al., 2017). However, widely used RL algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and REINFORCE Leave-One-Out (RLOO) (Ahmadian et al., 2024) suffer from a fundamental limitation: they are on-policy, meaning that data generation and policy updates occur sequentially. This dependence creates bottlenecks that reduce resource utilization. Moreover, the benefit of scaling on-policy data generation may be limited (Hou et al., 2024).
We introduce Trajectory Balance with Asynchrony (TBA), a distributed RL framework designed to efficiently and scalably leverage compute for LLM post-training. TBA decouples data generation from policy updates using an off-policy training objective based on Trajectory Balance (TB) (Malkin et al., 2022a). The framework uses multiple searcher nodes that independently generate diverse trajectories and store them in a central replay buffer, while a single trainer node asynchronously samples from this buffer to update the policy. By removing the dependency between data generation and training, TBA ensures high resource utilization and facilitates scalable search. By computing model updates with large, efficiently-generated off-policy response sets, the TB objective leverages response scaling to produce better models than popular RL baseline methods.
TBA offers three key advantages over existing RL-based LLM post-training approaches: 1) Asynchrony enables massive parallelization, significantly reducing training wall-clock time as shown in Figures 1 and 3. 2) Diverse off-policy sampling from a replay buffer improves exploration and prevents mode collapse. 3) Scalable search capabilities make TBA particularly effective in sparse reward settings, such as automated red-teaming.
We validate TBA on mathematical reasoning, preference-tuning, and automated red-teaming. Our results show that TBA achieves performance comparable to or better than existing methods while significantly improving training speed. Our key contributions are summarized as follows:
  • We introduce TBA, a novel distributed RL framework for LLM post-training.
  • We decouple data generation and policy updates, improving training speed and scalability.
  • We demonstrate the effectiveness of trajectory balance for LLM post-training, exploiting its ability to efficiently leverage large-scale off-policy data.
  • We demonstrate significant speedups (4x or more) in RL for mathematical reasoning, preference-tuning, and automated red-teaming.
By enabling high-quality and fast off-policy post-training, TBA contributes to the broader goal of scalable and effective LLM alignment, ensuring that large models can be refined more efficiently for real-world deployment.
2 Related Work
RL fine-tuning of language models
Reinforcement Learning (RL) has been an integral component for training LLMs (Google Gemini Team, 2024; OpenAI, 2023). In particular, RL has became the de facto approach for aligning language models with human preferences Christiano et al. (2017); Ziegler et al. (2019); Stiennon et al. (2020); Ouyang et al. (2022). Much of this work relies on Proximal Policy Optimization (PPO) (Schulman et al., 2017), an on-policy RL algorithm which has become a default choice for fine-tuning LLMs due to it’s strong performance across different setups. Aside from PPO, other on-policy objectives such as REINFORCE (Ahmadian et al., 2024) and variants like GRPO (Shao et al., 2024) and VinePPO (Kazemnejad et al., 2024) have also been studied in the context of language models.
An alternative to PPO-based fine-tuning is rejection sampling fine-tuning, inspired by the best-of-$n$ inference approach proposed by Nakano et al. (2021). Recent work by Dong et al. (2023); Gulcehre et al. (2023), and Wang et al. (2024) extends this concept by generating $n$ candidate responses for each prompt, ranking them with a learned reward function, and fine-tuning the model based on the highest-ranking responses. On the other hand, direct preference learning approaches (Rafailov et al., 2023; Azar et al., 2024; Tang et al., 2024) skip reward modeling entirely and train language models to directly optimize responses under a preference model. Finally, Hu et al. (2024) introduced GFlowNet fine-tuning, leveraging off-policy GFlowNet algorithms for fine-tuning language models, which we build upon in this work.
Asynchronous distributed RL
Distributed RL spreads actors/searchers, learners, and environments across a collection of computing resources. Asynchronous distributed RL does this such that searcher and trainer processes do not necessarily share the same weights, which can significantly improve training speed (Nair et al., 2015; Wang et al., 2025) and facilitate RL in complex, high-dimensional domains (Hessel et al., 2021; Huang et al., 2023; Horgan et al., 2018).
A foundational method in this area is Asynchronous Advantage Actor-Critic (A3C) (Mnih, 2016). In A3C, multiple parallel workers asynchronously interact with the environment and communicate gradients to a central node. Our approach to async distributed RL more closely resembles the Importance-Weighted Actor-Learner Architecture (IMPALA) method (Espeholt et al., 2018), which communicates experience trajectories (state, action, and reward tuples) to the central node.
Automated red-teaming
Through adversarial interactions, LLM red-teaming clarifies the robustness and risks of a target LLM. Automating the generation of these adversarial scenarios, automated red-teaming frameworks can help uncover vulnerabilities, biases, and unintended behaviors in models more quickly, enabling preemptive mitigation before deployment (Wei et al., 2023; Ji et al., 2024).
Perez et al. (2022) proposed training language models using RL to discover prompts that elicit harmful responses from some target LLM. Standard RL approaches, however, are susceptible to mode collapse and fail to achieve the attack diversity that is critical for successful red-teaming. Hong et al. (2024) introduced a curiosity bonus to encourage generation of diverse red-teaming prompts, whereas Samvelyan et al. (2024) proposed sampling an attack prompt from a pool and iteratively mutating the prompt with auxiliary LLMs. Lee et al. (2025) proposed using GFlowNet fine-tuning followed by MLE smoothing to generate diverse, transferable, and effective prompts. Our red-teaming experiments augment their TB objective optimization with our distributed asynchronous framework.
3 Preliminaries
KL regularized RL as probabilistic inference
We study the problem of fine-tuning a pretrained language model $\pi_{\text{ref}}$ with a reward model $r_{\phi}$. For mathematical reasoning (Cobbe et al., 2021) our reward model simply computes a response’s correctness, while our preference-tuning for alignment (Ziegler et al., 2019) and red teaming (Perez et al., 2022) experiments use reward models optimized with human preference data. Notably, reward maximizing RL with learned reward functions is susceptible to spurious modes of the reward (Skalse et al., 2022; Pan et al., 2022; Gao et al., 2023), resulting in poor performance and low diversity in responses. This is addressed by constraining the fine-tuned model to be close to the initial model in terms of the KL divergence:
$\begin{split}\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\mathbb{E}_{\mathbf{y}\sim\pi(\mathbf{y}\mid\mathbf{x})}[r_{\phi}(\mathbf{y};\mathbf{x})]\\ -\beta\mathbb{D}_{\text{KL}}(\pi(.\mid\mathbf{x})||\pi_{\text{ref}}(.\mid\mathbf{x}))].\end{split}$ (1)
Online RL algorithms such as PPO (Schulman et al., 2017; Stiennon et al., 2020) and REINFORCE (Williams, 1992; Kool et al., 2019; Ahmadian et al., 2024) can be used to optimize the model (which is the policy), whereas offline objectives such as DPO (Rafailov et al., 2023) can be used to train the model directly on the preference data and achieve the same optimal policy asymptotically.
Eq. 1 can be interpreted from a probabilistic perspective as a Bayesian posterior inference problem (Korbak et al., 2022). The optimal policy for Eq. 1 is given as:
$\pi^{*}(\mathbf{y}\mid\mathbf{x})\propto\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})\exp(\beta^{-1}r_{\phi}(\mathbf{y};\mathbf{x})).$ (2)
Approaches such as Gibbs sampling can be used to produce samples from the optimal policy without any fine-tuning (Xiong et al., 2024). These MCMC approaches can be very expensive at inference time and intractable for long sequences. Within the probabilistic interpretation, on-policy RL to maximize Eq. 1 is equivalent to amortized variational inference to minimize the reverse KL with respect to the posterior density (Korbak et al., 2022). However, reverse KL optimization is susceptible to mode collapse and requires on-policy samples. The reliance on on-policy samples can limit the scalabality as we cannot use replay buffers that can be populated in parallel at scale, since new updates of the policy invalidate the on-policy nature of the older replay buffer examples. In practice this means that the policy can get stuck in suboptimal solutions and stop exploring, or may obtain high reward at the cost of diversity. This motivates us to consider an alternative off-policy amortized variational inference to efficiently leverage scalable computational resources for flexible exploration.
GFlowNets
Generative Flow Networks (GFlowNets; Bengio et al., 2021; 2023) are a framework for off-policy training of hierarchical generative models to sample proportional to a given unnormalized density (reward) function. GFlowNets frame probabilistic inference as a sequential decision-making problem, learning a policy to construct the objects (e.g. sequences) by putting together building blocks (e.g. tokens) and optimizing consistency-based objectives. GFlowNet objectives have been used for fine-tuning autoregressive (Hu et al., 2024; Lee et al., 2025) as well as discrete diffusion language models (Venkatraman et al., 2024).
To fine-tune a language model to sample from Eq. 2, we can set as a reward $R(\mathbf{y};\mathbf{x})=\pi_{\text{ref}}(\mathbf{y}\mid\mathbf{x})\exp(\beta^{-1}r_{\phi}(\mathbf{y};\mathbf{x}))$. Following Lee et al. (2025), we use the trajectory balance objective (Malkin et al., 2022b) for training the language model policy $\pi_{\theta}$, which is defined over a response $\mathbf{y}$ as
$\mathcal{L}_{\text{TB}}(\mathbf{y},\mathbf{x};\theta)=\left(\log\frac{Z(\mathbf{x})\pi_{\theta}(\mathbf{y}\mid\mathbf{x})}{R(\mathbf{y};\mathbf{x})}\right)^{2}.$ (3)
$Z(\mathbf{x})$ is a positive scalar function of the query $\mathbf{x}$, and the response y is a sequence of tokens. When $\mathcal{L}_{\text{TB}}$ is minimized, $Z(\mathbf{x})$ is the partition function of the posterior (i.e. $Z(\mathbf{x})=\sum_{\mathbf{y}}R(\mathbf{y};\mathbf{x})$). Instead of training a value network for $Z$, we use the VarGrad variant of trajectory balance which replaces a learned $Z$ with a batch estimate  (Richter et al., 2020; Nüsken & Richter, 2021; Zhang et al., 2023; Sendera et al., 2024; Venkatraman et al., 2024).
Given $K$ responses $\{\mathbf{y}^{(i,j)}\}_{j=1}^{K}$ for a query $\mathbf{x}^{(i)}$, a batch estimate of $Z$ can be computed as follows
$\begin{split}\log\hat{Z}(\mathbf{x}^{(i)})=\frac{1}{K}\sum_{j=1}^{K}\bigg{(}\sum_{t=1}^{T}\log\pi_{\text{ref}}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})-\log\pi_{\theta}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})+\frac{1}{\beta}r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\bigg{)}.\end{split}$ (4)
The estimate $\hat{Z}$ can be plugged into Eq. 3 for a batch $\mathbf{B}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i,j)})_{j=1}^{K}\}_{i=1}^{B}$ to get
$\begin{split}\mathcal{L}_{\text{TB}}^{\text{VarGrad}}(\mathbf{B};\theta)=\frac{1}{BK}\sum_{i=1,j=1}^{i=B,j=K}\Bigg{(}\log\hat{Z}(\mathbf{x}^{(i)})+\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})\\ -\log\pi_{\text{ref}}(y_{t}^{(i,j)}\mid y_{<t}^{(i,j)},\mathbf{x}^{(i)})-\frac{1}{\beta}r(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\Bigg{)}^{2}.\end{split}$ (5)
An important property of the trajectory balance is that it is off-policy. During training, $\mathbf{y}$ can be sampled from any distribution with full support (Bengio et al., 2021). This enables the use of various exploration strategies (Rector-Brooks et al., 2023; Kim et al., 2024) as well as the use of replay buffers (Shen et al., 2023; Vemgal et al., 2023). In the context of fine-tuning language models, this off-policy nature of the objective makes it a natural choice for large-scale distributed training.
Figure 2: Fast, scalable LLM post-training with TBA. Continuously (solid lines), multiple Searcher nodes (left) collect trajectories, while a Trainer node (right) samples from a replay buffer to train the policy off-policy. Periodically (dashed lines), updated policy weights are sent to Searcher nodes, and new trajectories are added to the Trainer node’s buffer. This avoids bottlenecks at any given node, which can be 1 or more GPUs, keeping resource utilization high.
4 TBA: Fast, Scalable LLM Post-Training
TBA is an asynchronous distributed RL framework for post-training language models (see Figure 2 for a visualization). TBA uses the off-policy trajectory balance objective (Equation 5) to efficiently leverage scaled data generation, and it decouples data generation from model updates to ensure high resource utilization. TBA has two key components: a single Trainer node and one or more Searcher nodes that collect off-policy data into a shared replay buffer $\mathcal{D}_{\mathrm{global}}$.
TBA can be implemented by, e.g., modifying the RLOO trainer class of the Hugging Face TRL library (von Werra et al., 2020) to use the TB objective, creating a dictionary to hold trajectories, and assigning training and search to distinct nodes. Here, a node is a computational resource sufficient to perform all the operations needed for training or search – it can be a collection of multiple GPUs or even a subset of 1 GPU. In our experiments, a node is always 1 GPU; e.g., given 16 GPUs, we would have 15 Searcher nodes and 1 Trainer node operating.
Separation of the Searcher and Trainer is highly desirable even in a 2 node (e.g. 2 GPU) cluster because LLM policy rollouts are costly sequential decoding procedures, and training requires only parallel likelihood evaluation of an entire sequence through a single forward pass. Thus, given an objective that can tolerate off-policy asynchronous online training, massive wall clock time speedups can be realized by continuously running training on 1 node without pausing for rollout generation.
4.1 Scaling data collection with Searcher nodes
In TBA, the Searcher nodes each carry a local delayed copy $\pi_{\theta^{\prime}}$ of the Trainer policy $\pi_{\theta}$. To produce policy rollouts, queries $\mathbf{x}$ are sampled from a dataset, and the local policy generates a batch of $K$ responses $\mathbf{y}\sim\pi_{\theta^{\prime}}(\mathbf{y}|\mathbf{x})$ that are evaluated with the reward model $r_{\phi}(\mathbf{y};\mathbf{x})$. Like Noukhovitch et al. (2025), we use vLLM (Kwon et al., 2023) for faster generation. The $(\mathbf{x},\mathbf{y},r_{\phi}(\mathbf{y};\mathbf{x}))$ tuples are stashed in the Searcher’s local replay buffer $\mathcal{D}_{\text{local}}$ (note that $\mathbf{x}$ can instead be stored as a dictionary key to save space). We also add to the stored tuple the step of the trainer when syncing last occurred, giving us a notion of how off-policy the generated data is – this can later be used by the Trainer to prioritize sampling from more recent generations for relatively “on-policy” updates.
Periodically, search and training pause to pull each Searcher node’s local replay buffer $\mathcal{D}_{\text{local}}$ into the global replay buffer $\mathcal{D}_{\mathrm{global}}$, and to update the searcher’s local policy with the trainer’s version. The global buffer maintains a list of all generated responses and rewards for each query.
A key motivation for scaling up data generation through a large number of Searcher nodes is improving exploration. For example, we generate $S>K$ samples for a given query, even when only updating the model using $K$ samples per query – this can mitigate the lack of diversity caused by the fact that $K$ independent model rollouts are not guaranteed to produce $K$ unique sequences (when duplicates are encountered, we keep the most recently generated version). Relatedly, future work could apply simple off-policy inference techniques in the Searcher nodes such as randomly sampling the softmax temperature, or using alternative decoding techniques like beam search. We expect such approaches to particularly aid solution discovery in sparse reward settings.
4.2 Asynchronous updates with Trainer
The Trainer uses off-policy trajectory balance (Eq. 5) to train the policy on the global replay buffer $\mathcal{D}_{\mathrm{global}}$. We sample a batch of $B$ queries, each with $K$ responses and corresponding rewards:
$\{\mathbf{x}^{(i)},\mathbf{y}^{(i,j)},r_{\phi}(\mathbf{y}^{(i,j)};\mathbf{x}^{(i)})\}_{i=1,j=1}^{i=B,j=K}\sim\mathcal{D}_{\text{global}}.$
We then compute the loss in Eq. 5 and use it to update the policy. We sample with replacement if fewer than $K$ unique samples exist for a given query.
A critical design choice is the strategy for sampling from the replay buffer $\mathcal{D}_{\mathrm{global}}$. The most naive approach is uniform sampling over queries, then uniform sampling over samples associated with the selected query, which may not be optimal if high-reward samples are sparse. Reward prioritization can address this by tilting the sampling distribution toward high-reward sequences; however, focusing solely on high-reward data can lead to mode collapse and reduce policy diversity.
To balance between these concerns, we alternate between two sampling strategies: one prioritizing recency – i.e., whether the trajectory was added to the buffer in the most recent sync step – and another prioritizing rewards. When prioritizing rewards, we consider both a softmax of the reward value (to encourage sampling high reward responses) and a uniform distribution (to encourage sampling high and low reward responses equally). We randomly switch between prioritizing rewards and recency for each query in a batch, with the fraction of queries allocated to each strategy treated as a tunable hyperparameter $m$, which we study in Section 5.4.
5 Experiments
We evaluate the effectiveness of TBA in three common LLM post-training RL pipelines. Post-training RL for enhancing LLM capabilities—particularly for agentic and reasoning tasks—is a critical but nascent area where baseline approaches require many hours or days. These conventional methods often rely on costly-to-generate on-policy data and thus inefficiently leverage available computing resources. Notably, this inefficiency can become particularly harmful when scaling to larger distributed systems, which may be a necessity for domains with sparse rewards that demand increased sampling. Broadly, we find TBA is a highly-performant, fast, and scalable solution.
5.1 Tasks
  • Mathematical reasoning (MR): We study the GSM8K task which consists of grade-school level math problems and a binary reward based on exact match for the correct final answer (Cobbe et al., 2021). We adopt the setup from Kazemnejad et al. (2024); Noukhovitch et al. (2025), using an SFTed RhoMath-1B (Lin et al., 2024) model as a base for RL post-training.
  • Preference fine-tuning (PFT): We consider the task of fine-tuning a language model with a reward function learned from human preference data. Specifically, we study the TL;DR summarization task where the goal is to write short summaries for reddit posts (Stiennon et al., 2020). Following Noukhovitch et al. (2025), we consider Pythia (Biderman et al., 2023) as the base model for the policy and the reward models, using the SFTed versions used by Noukhovitch et al. (2025).
  • Red-teaming (RT): We investigate automated red-teaming, another critical step for LLM post-training. The goal is to discover prompts that elicit harmful responses from a target model, as measured by a toxicity classifier. We follow the setup from Lee et al. (2025), applying the same models: our smaller-scale experiments use GPT-2 (Radford et al., 2019) as an attacker model, GPT-2 (instruction-tuned) as a victim model, and a RoBERTa-based toxicity classifier (Vidgen et al., 2021); our larger-scale experiments use Llama-3.2-1B (Meta AI Llama Team, 2024) as an attacker model, Llama-3.1-8B-Instruct as a victim model, and LlamaGuard-3-8B (Meta AI Llama Team, 2024) for measuring toxicity, averaging over multiple responses from the victim model.
For all tasks, we study solely the RL post-training components of baselines’ workflows. In particular, we start with SFTed checkpoints, then perform RL. Notably, Lee et al. (2025) also include a maximum likelihood estimation training phase after RL that uses the buffer produced during RL, which boosts performance but is not investigated here. Using these baselines’ codes, we follow their setup and hyperparameters, with deviations noted in Appendix A. For MR and PFT, we implement TBA by augmenting the RLOO trainer of Noukhovitch et al. (2025) with our distributed asynchronous RL framework and the TB objective (Equation 5). For RT, we implement TBA by augmenting the trajectory balance trainer of Lee et al. (2025) with our distributed asynchronous RL framework.
5.2 Metrics and baselines
MR Metrics. We follow the evaluation setup of Noukhovitch et al. (2025), computing the pass@1 on the GSM8K (Cobbe et al., 2021) test dataset with greedy decoding.
MR Baselines. Prior work (Kazemnejad et al., 2024; Noukhovitch et al., 2025) applies RL post-training with the GSM8K training set to the SFTed RhoMath-1B (Lin et al., 2024) model, which initially obtains 40.3% accuracy on the test set. We compare TBA post-training to the methods used in these prior works: VinePPO (Kazemnejad et al., 2024), Online-DPO (Guo et al., 2024), PPO (Schulman et al., 2017), and RLOO (Ahmadian et al., 2024).
PFT Metrics. We follow the evaluation setup of Noukhovitch et al. (2025), using the win-rate under a 6.7B “gold” reward model (Huang et al., 2024) as the primary metric. We additionally report approximate KL distance—approximated by perplexity to match the evaluation of Noukhovitch et al. (2025)—between the learned policy and the reference policy.
PFT Baselines. Following Noukhovitch et al. (2025), we compare TBA with: Online-DPO (Guo et al., 2024), PPO (Schulman et al., 2017), and RLOO (Ahmadian et al., 2024).
RT Metrics. We follow Lee et al. (2025), measuring the attack success rate on 1024 sampled prompts for a victim model. We also measure the diversity of these test-time generated prompts by computing the average pairwise cosine distance.
RT Baselines. We compare against: SFT, PPO (Schulman et al., 2017) (with the novelty reward from Lee et al. (2025)), REINFORCE (Williams, 1992; Sutton et al., 1999), RLOO (Ahmadi & Mahmudi, 2023), Online DPO (Rafailov et al., 2023), GFlowNet (one-actor TB) (Lee et al., 2025).
Figure 3: TBA scales search and improves RL efficiency on the TL;DR summarization task. All plotted points use 4xA100 GPUs, but TBA allocates 3 GPUs to search, and Online DPO allocates 1 GPU to search. TBA produces large-scale off-policy data that its trajectory balance objective can leverage, creating massive efficiency benefits. Online DPO baselines taken from Noukhovitch et al. (2025). Dashed and solid lines use 256 and 425 updates, respectively. Appendix A has details.
Figure 4: TBA defines a new KL vs. win-rate Pareto frontier for the TL;DR summarization task. The baseline “Online DPO” frontier is created by increasing the degree of off-policyness, starting from on-policy Online DPO, results from (Noukhovitch et al., 2025). The TBA frontier is created by altering the training steps, searcher count, and KL annealing schedule as described in Appendix A.
Table 1: TBA Pareto dominates baseline on-policy algorithms in the PFT task (Pythia 410M). Baseline results taken from Noukhovitch et al. (2025). Please see Appendix A for details.
Method Perplexity/KL $\downarrow$ Win Rate $\uparrow$
Online DPO 1.13 0.85
PPO 1.14 0.86
RLOO 1.13 0.82
TBA (Ours) 1.13 0.86
5.3 TBA redefines efficiency-performance Pareto frontiers
We hypothesize that, by mixing asynchronous RL with the trajectory balance objective for learning from off-policy data, TBA can (a) reduce resource waste and training time and (b) improve performance via scaled generation and ingestion of diverse responses. To test this, we primarily consider compute-matched settings where all methods have access to the same number of GPUs, though we also evaluate TBA with resource scaling. Even in our compute-matched experiments, TBA generates responses relatively rapidly by running asynchronous search on all but one of the available GPUs. Training happens quickly with TBA because it isn’t bottlenecked by on-policy generation, and TBA’s rapid response generation ensures training on diverse and previously unseen (though off-policy) data.
When tested on established RL problems, an alternative possibility is that TBA will underperform due to its departures from conventional approaches: it is asynchronous, off-policy, reliant on the less-common trajectory balance objective, and makes updates using many responses per query. [^1] We compute the TBA loss with more samples per query than analogous approaches like RLOO (e.g., 20 samples versus 4) to reduce the variance of the gradient of the TB objective estimate (Equation 5) that we optimize. We use fewer queries per batch to keep the batch size small despite this response scaling. Indeed, Noukhovitch et al. (2025) contemporaneously suggests potential limitations to asynchronous, off-policy RL for LLMs, finding that increasing off-policyness can harm performance metrics like win-rate or exacerbate policy deviations from the reference model. Further, Hou et al. (2024) found limited benefits to scaling response generation from 4 to 8 (or 16) responses when using PPO.
We study this question by computing Pareto frontiers for MR (Figure 1), for PFT (Figures 3 and 4 and Table 1), and for RT (Figure 5 and Table 5.4). See Appendix A for experimental details. Notably, TBA produces results on or beyond the Pareto frontiers of all three tasks at multiple model scales, consistent with our hypothesis that TBA can efficiently train LLMs on off-policy data.
Speed is vastly improved with TBA training, which proceeds entirely asynchronously without training-bound or generation-bound processes – the only non-training time occurs briefly every $k$ steps. In the compute-matched MR experiments (Figure 1), TBA speeds up the training of the only method with comparable performance (VinePPO) by nearly 50×50$\times$50 ×, while improving accuracy by $1.8$% and speed by 1.5×1.5$\times$1.5 × relative to the speed-optimized asynchronous DPO baseline (Noukhovitch et al., 2025). In the compute-matched PFT experiments (Figure 3), TBA produces 5$\approx$ 5$\times$ 5 × speedups over speed-optimized asynchronous DPO baselines (Noukhovitch et al., 2025). In the non-compute-matched automated red-teaming experiments (Section 5.4), TBA gives 7$\approx$ 7$\times$ 7 × speedups for GPT-2 and Llama 3.2 1B compared to the non-distributed, synchronous GFlowNet baseline (Lee et al., 2025). these results suggest that TBA is an effective, parallel, and scalable search framework for distributed learning, offering substantial speed-ups while remaining competitive with leading approaches.
5.4 Does off-policyness hurt performance?
The results in the prior section are perhaps surprising given evidence of off-policyness’s harmfulness to RL post-training (Noukhovitch et al., 2025). Thus, we investigate the effect of off-policyness with TBA by studying its related hyperparameters, which we first review here. The fraction $m$ controls how often sampling attempts to approximate an “on-policy” distribution. When $m=1$, training occurs exclusively with samples added in the most recent sync step, while $m=0$ corresponds to selecting data without regard for how recently it was generated (e.g., using reward weighting). Additionally, recall that parameters and buffer data are shared between searchers and trainers every $k$ training steps: since TBA runs exploration and training in parallel, TBA trains off-policy even when $m=1$ and $k=1$. Specifically, with probability $m$, TBA selects a sample from the central replay buffer that is at most $12k-$ 1 updates off-policy. With probability $1-m$, TBA samples data produced by the model at any point, which can be as off-policy as the number of training steps minus $1$.
To understand the effect of increasing off-policyness on TBA, we test the effect of modifying $m$ for the MR (see Figure 6) and PFT tasks. For PFT, we train Pythia-410M on the TL;DR dataset with three values of $m$ ($0.4,0.5,0.6$) keeping all other parameters constant. We found that win rate fluctuated, with $m=0.4$ corresponding to the lowest win rate ($0.67$), and $m=0.5$ and $m=0.6$ attaining higher win rates of $0.82$ and $0.8$, respectively. These results show that higher values of $m$ generally lead to a higher win rate, reinforcing the idea that on-policy updates are in fact the most effective. However, for reasonably high values of $m$, incorporating more off-policy data does not significantly degrade performance and, in some cases, may even provides benefits. Regardless, our findings for both MR and PFT further support the idea that we can perform massively distributed training that works well with off-policy updates, as long as recent samples are thrown into the mix.
Importantly, the choice of reinforcement learning (RL) algorithm is crucial in determining how effectively we can leverage off-policy updates. TB is a fully off-policy compatible algorithm, and as shown in Figure 4, it significantly outperforms asynchronous Online DPO, even in the latter’s most on-policy setting. Noukhovitch et al. (2025) identified Online DPO as the best-performing off-policy algorithm in their experiments, making it particularly insightful that TB improves upon this.
Figure 5: TBA reaches the RT diversity-toxicity Pareto frontier and improves as search is scaled. (Left) On the GPT-2 automated red-teaming task of Lee et al. (2025), TBA produces results on the diversity vs. toxicity Pareto frontier in less training time. Baselines taken from Lee et al. (2025). (Right) Each searcher uses one V100 GPU for generating attacks. We report means and standard errors from multiple runs of the automated red-teaming task with GPT-2 at each searcher/GPU count.
Table 2: TBA speeds up the wall-clock time required to reach the Pareto frontier for the red-teaming task. The GFlowNet performances are taken from Lee et al. (2025), while the training speeds are computed by us with their code. With the GPT-2 models, TBA performance improves with searcher count. With the Llama models, we trade attack toxicity for attack diversity by scaling the TBA buffer’s maximum size from 130,000 to 150,000 samples*, retaining more off-policy data.
{widetable}
lllrlll Attacker / Training Method Hardware Time (h) Speedup Cosine % Toxic
Victim Model Distance Prompts
GPT-2 / GPT-2 + SFT
GFlowNet - Sync 1$\times$×V100 11.9 1x 0.65 96.6
TBA (Ours) 4$\times$×V100 1.7 7x 0.42 94.5
TBA (Ours) 16$\times$×V100 2.5 4.8x 0.45 96
TBA (Ours) 64$\times$×V100 2.9 4.1x 0.49 97.6
Llama 3.2 1B / Llama 3.1 8B - Instruct GFlowNet - Sync 1$\times$×A100 37.4 1x 0.32 100.0
TBA (Ours) 8$\times$×A100 5.7 6.6x 0.35 98.1
TBA (Ours)* 8$\times$×A100 5.7 6.6x 0.37 94.8
5.5 Discovery of high-reward samples via scaling search
Beyond improvements at a given level of compute, we also observe improvements when we scale the amount of total compute, showing TBA’s promise for large-scale distributed RL. In our asynchronous setup, we find that adding more searchers consistently improves the attack success rate and diversity for RT (see Figure 5, right). This improvement likely stems from having more searchers exploring different regions of the solution space simultaneously, enabling more effective discovery of high-reward samples. Moreover, asynchronous updates introduce opportunities for beneficial randomness, and thus potential expansion of the search coverage in the combinatorial space of language. Interestingly, we also find some evidence for scaling’s helpfulness in PFT (see Figure 7).
6 Discussion
In this work, we introduced TBA, a novel post-training method for large language models (LLMs) that combined an off-policy reinforcement learning (RL) objective with distributed asynchronous search. By decoupling searcher and trainer nodes, our approach enabled efficient distributed training and avoided bottlenecks, leading to significant performance gains in post-training tasks such as mathematical reasoning, automated red-teaming, and RLHF. We expect that our highly parallelizable and performant framework for RL post-training can be extended to other valuable tasks, including self-improvement training (Zelikman et al., 2022; Hu et al., 2024; Gulcehre et al., 2023), search-based reasoning (Wan et al., 2024), as well as the emerging paradigm of training LLMs to reason with RL (Guo et al., 2025).
Towards a multi-agent search system
Our approach can potentially be extended to a multi-agent search system aimed at identifying multiple diverse modes within the large combinatorial space of language. Our current searchers do not explicitly target different regions, but we do show that multiple distributed searchers can enhance exploration and populate a global replay buffer with varied experiences. By evolving this concept into an explicit multi-agent system, we can encourage each agent to explore distinct regions in a divide-and-conquer manner. This cooperative strategy allows agents to effectively identify and report local modes to a centralized replay buffer. The off-policy post-training agent can then learn from the diverse samples stored in this centralized buffer, benefiting from a wider range of experiences. We consider this a promising future direction.
Improving local credit assignment
The trajectory balance objective can suffer from high gradient variance as it operates on the trajectory level. We addressed this by sampling more responses per query. Future work can leverage learning partial energy functions (Madan et al., 2023; Zhang et al., 2025) to balance bias and variance during policy updates.
Acknowledgements
The authors acknowledge funding from CIFAR, NSERC, IVADO, and Samsung.
Prepared by LLNL under Contract DE-AC52-07NA27344 and supported by the LLNL-LDRD Program under Project No. 24-ERD-058 (LLNL-CONF-2003261). This manuscript has been authored by Lawrence Livermore National Security, LLC under Contract No. DE-AC52-07NA27344 with the U.S. Department of Energy. The United States Government retains, and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC awards ASCR-ERCAP0032802 and ASCR-ERCAP0032812.
Appendix A Experimental Details
As discussed in Sections 4 and 5.4, TBA introduces new hyperparameters. Most notably, these include (1) the sync period $k$, which is the number of training steps between two successive model-buffer synchronizations; and (2) the most-on-policy probability $m$, which is the probability of sampling the data that is most on-policy – i.e., the samples that were added to the buffer during the most recent synchronization. When listing hyperparameter values, we clarify those that are specific to TBA and include references to their discussion/visualization in the text.
For MR and PFT, we implement TBA by building on the RLOO trainer class of the Hugging Face TRL library (von Werra et al., 2020). For MR and PFT, we make modifications to the baseline hyperparameters as shown in Table 3 and Table 4, respectively.
Our RT implementation of TBA built on the TB trainer used in Lee et al. (2025), and we largely utilize their code’s default hyperparameters, with differences noted in Appendix A.3.
A.1 GSM8K Mathematical Reasoning (MR)
All baselines and TBA results use the same starting point, a RhoMath-1B (Lin et al., 2024) model SFTed on GSM8K training data by Kazemnejad et al. (2024) – “realtreetune/rho-1b-sft-GSM8K” on Hugging Face. The baseline model achieves $40.3$% test accuracy. For the VinePPO baseline, training time is estimated using their reported 380 seconds per training step and 650 steps for GSM8K training (Kazemnejad et al., 2024).
Hyperparameter Value Reference
Model Rho-1B SFT on GSM8K
Learning Rate $1\times 10^{-5}$
Learning Rate Schedule Warmup Stable Decay
Learning Rate Warmup Steps 50
Learning Rate Stable Steps 450
Learning Rate Decay Steps 500
Generation Temperature 0.7
Max Prompt Token Length 512
Response Length 512
Number of Prompts per Batch 7
Number of Completions per Prompt 20 $K$ in Section 4.1
Batch Size (effective) 140
Number of Training Steps 1000
Total Prompts Seen 7000
Total Episodes 140000
TBA-specific hyperparameters
Beta (KL coefficient) Initial Value 0.012 $\beta$ in Equation 5
Beta Final Value 0.004 $\beta$ in Equation 5
Beta Linear Decay Schedule End Step 500 $\beta$ in Equation 5
Number of Samples per Prompt 24 $S$ in Section 4.1
Most-On-Policy Sampling Probability 0.95 $m$ in Section 5.4
Sync Period 2 $k$ in Section 5.4, Figure 2
Number of Searchers 3 Searchers in Figure 2
Number of Trainers 1 Trainer in Figure 2
Reward-Based Sampling Prioritization Uniform Section 4.2
Initial Completions in Buffer 500 Buffer at Step 0 in Figure 2
Table 3: TBA Training Hyperparameters for the GSM8K MR task.
In Figure 1, TBA uses the settings in Table 3 with a few modifications to shorten training time by an additional $30$%, down to 82 minutes on 4xA100 GPUs. In particular, we observed in initial testing (see Appendix B) that using the hyperparameters in Table 3 led to no improvement in performance for the final 300 steps. Thus, we shrank the training duration from 1000 steps to 700 (98000 episodes with batch size 140). Additionally, we made the following modifications to attempt to reduce the variance of the shortened run: 350 stable learning rate steps (down from 450), 0.014 Beta Initial Value (up from 0.012). We ran this experiment three times – obtaining performances of $55.8$%, $53.9$%, and $54.1$% – and reported the mean accuracy $54.6$% in Figure 1.
Limitations and future work
The 700-step result we show in Figure 1 has standard error $0.6$%, which is a little more variance than what we observe in the original 1000 step setup (the blue line in the bottom left plot of Figure 6 shows the mean and standard errors for the 1000 step runs). Future work could further explore variance reduction approaches/hyperparameters for shorter runs (and for TBA/RL more generally).
One way to deal with variance is to choose a checkpoint based on a held out validation set, a strategy used to produce the VinePPO result (Kazemnejad et al., 2024). We do not use this approach but note that our results would likely benefit significantly ($\approx 1$%) from it. In particular, each of our runs tends to produce (at some point during training) a higher performance than the performance at the final step – this is expected if you consider that the model performance varies around the average value it converges to towards the end of training (e.g., see again the blue line in the bottom left plot of Figure 6). Despite its being lower than the maximum performance achieved at any point during training, we report this final step performance, which is what a user of TBA could expect to obtain without applying techniques like early-stopping.
Hyperparameter Value Reference
Model Pythia SFTed on TL;DR
Learning Rate $3\times 10^{-6}$
Learning Rate Schedule Linear
Generation Temperature 0.7
Max Token Length 1024
Max Prompt Token Length 512
Response Length 128
Number of Prompts per Batch 8
Number of Completions per Prompt 20 $K$ in Section 4.1
Batch Size (effective) 160
Number of Training Steps 256
Total Prompts Seen 2048
Total Episodes 40960
TBA-specific hyperparameters
Beta (KL coefficient) Initial Value 1 $\beta$ in Equation 5
Beta Final Value 0.05 $\beta$ in Equation 5
Beta Linear Decay Schedule End Step See caption $\beta$ in Equation 5
Number of Samples per Prompt 20 $S$ in Section 4.1
Most-On-Policy Sampling Probability 0.5 $m$ in Section 5.4
Sync Period 10 $k$ in Section 5.4, Figure 2
Number of Searchers 3 Searchers in Figure 2
Number of Trainers 1 Trainer in Figure 2
Reward-Based Sampling Prioritization Softmax of Score Section 4.2
Initial Completions in Buffer 10000 Buffer at Step 0 in Figure 2
Table 4: TBA Training Hyperparameters for the TL;DR PFT task. For the PFT task, we accelerate the decay of Beta. In particular, the Beta Linear Decay Schedule End Step is set to be half the number of training steps, but we abruptly end this decay and set Beta to its final value at one eighth the number of training steps (e.g., step 32 for 256 steps). This has the effect of trading off KL/perplexity, which we found to be relatively low with our TBA setup, for win rate.
A.2 TL;DR Preference Fine Tuning (PFT)
For PFT in Figure 3, we use the settings in Table 4 as well as a longer-duration run with 425 updates.
For PFT in Figure 4, we create the TBA frontier by modifying the training steps, searcher count, and beta linear decay schedule shown in Table 4. We train for 256, 425, 625, 725, and 825 steps, and we search with 2, 3, 4, and 7 searcher nodes. We did not notice a significant pattern in performance when changing searcher count, but we found a tendency for higher KL/perplexity values and win-rates with more training steps (an expected tradeoff from optimizing the policy further). We noticed that the 2.8B model did not create a wide range of KL/perplexity values with these initial runs, so we also performed runs at that model size (with 825 steps, and with 2 and 4 searchers) using a less rapid beta linear decay schedule (reaching the beta final value at step 140 instead of step 104). This schedule change had the effect of reducing the KL/perplexity and win-rate (another expected tradeoff).
For PFT in Table 1, we use the settings in Table 4, except we train for 625 steps (100000 episodes) because we found use of more steps tended to improve win rate without a significant increase in perplexity (approximate KL). Additionally, we only use 2 searchers in this run. See Figure 7 for a depiction of the effects of step count and searcher count.
Limitations and future work
All of our PFT results were run in 32-bit precision and without DeepSpeed, which was used by baselines we compared against (Noukhovitch et al., 2025). Lower precision and such training acceleration packages could further improve the speedups we show. Relatedly, for our $2.8$B runs, we used gradient checkpointing to fit more examples into a micro batch, which led to slowdowns at this scale (i.e., we only have a $3.8$x speedup over the baseline in this setting). We leave the optimization of our framework with appropriate packages to future work. Finally, we used a large number (10000) of initial completions in the buffer, and future work should confirm that a smaller number (e.g. 1000) works equally well – note that a small number worked well for GSM8K.
A.3 Automated Red Teaming (RT)
Unlike our MR and PFT implementations, our RT implementation uses the code of Lee et al. (2025) as a baseline and thus does not follow the Hugging Face TRL trainer style. We discuss hyperparameters in the context of their trajectory balance trainer approach below. We adopt their code largely as it is, with the exception that our TBA implementation uses larger replay buffers than Lee et al. (2025) to accommodate our scaled search for trajectories. Additionally, unlike Lee et al. (2025), we remove the oldest samples when the buffers fill up as opposed to the lowest reward samples.
For the Llama results in Table 5.4, our hyperparameter choices largely follow those of Lee et al. (2025). We train for 5000 steps with batch size 128. For temperature sampling, we use a low and high of 0.7 and 2.0, respectively. We use a reward schedule horizon of 1000. The language model schedule end is 1.2, and its schedule’s horizon is 2000. We prioritize sampling based on reward. We use Beta 0.05. We use sync period ($k$) 10. We use 6 searchers. We use most-on-policy probability ($m$) 0.5 and 0.6 in combination with maximum buffer sizes 150000 and 130000, respectively. By having a smaller maximum buffer size and larger $m$, the latter setting is expected to focus on more recently generated data and prioritize reward over diversity, which is what we observe in Table 5.4.
For the GPT2 results in Figure 5 and Table 5.4, our hyperparameter choices again largely follow those of Lee et al. (2025). We train for 5000 steps with batch size 128. We use Beta 0.05. We use sync period ($k$) 10. We use most-on-policy probability ($m$) 0.5. We cap the maximum buffer size at 100000 in all experiments; we additionally prevent the buffer from growing past its size as of step 4000 to encourage focusing on more recent data (larger most on policy probabilities $m$ may provide a similar effect). We test searcher counts 2, 6, 14, 30, 62.
Limitations and future work
In Table 5.4, there is a slowdown as the number of searchers scales. We believe this is largely addressed by a newer version of our code that uses more efficient buffer communication, but we have not re-run these results yet to confirm this. In any case, developing more efficient TBA implementations is an interesting direction for future work given TBA’s ability to effectively leverage large-scale data generation.
Appendix B GSM8K Ablation Studies
Figure 6: GSM8K ablation studies. All experiments begin with the base hyperparameters listed in Table 3 and make the depicted modifications, except when studying the synchronization period $k$ in the top right plot (where we use Beta Final Value $0.005$ because $0.004$ led to instability with $k=4$). We report the mean and standard error from 2 runs of each configuration.
We adopted the hyperparameters for our GSM8K result in Figure 1 based on a series of trial experiments centered around the hyperparameters shown in Table 3. In Figure 6, we show the effects of changing what we found to be key hyperparameters. We note the following observations about TBA hyperparameter effects on GSM8K.
  1. 1.
    Unlike PFT, it was important for $m$ to be somewhat large for the GSM8K MR task (Figure 6, top left). Similarly, syncing more frequently was beneficial (Figure 6, top right). Together, these results suggest that GSM8K performance is more sensitive to off-policyness than performance on other tasks (e.g., PFT).
  2. 2.
    We found that the WSD schedule could add stability (Figure 6, bottom right).
  3. 3.
    Using smaller Beta Final Values tended to improve performance (Figure 6, bottom left), but training became unstable around $0.003$. This suggests a tradeoff between stability and accuracy for GSM8K that is mediated by the KL coefficient $\beta$.
  4. 4.
    A batch size of 140 did not provide significantly better or worse results than larger batch sizes in initial testing, but smaller batch sizes allow for faster training steps, motivating our use of 140.
Appendix C TL;DR Ablation Studies
Figure 7: TL;DR ablation studies. All experiments begin with the base hyperparameters listed in Table 4 then make the depicted modifications. More searcher nodes and more training steps tend to improve win rate at the cost of higher perplexity.
With TBA, we take many more optimization steps per hour of training than competing methods take. Accordingly, we sought to understand how taking more steps or using more searchers (to reach compute-parity with competing methods) affects performance. As shown in Figure 7, win rate tends to improve with increased compute through changes to these variables, though perplexity suffers as expected when we over-optimize the policy (in the case of step scaling). It is not clear why scaling the searcher count seems to have an effect similar to step scaling. However, there is a notable mechanism through which searcher scaling could have an effect: using more searchers should reduce the probability of selecting a training prompt that’s been seen before when sampling off-policy (because scaling the searchers scales the number of unique prompts with completions added to the buffer). [^2] Notably, the effect of scaling search is different in the case of RT, where there is a single prompt and scaling searchers makes it more likely that higher-reward samples are found and trained on. Here, with PFT, scaling search means that more prompts per second have a set of completions generated for them, but any given prompt is not expected to have higher-reward samples as a result of scaling the searcher count – this is an implementation choice and could be changed to gain the effect that scaling searchers has with RT, however. However, the effect size is small and inconsistent enough to suggest this searcher-scaling trend (in the context of PFT) needs further investigation before it’s confirmed.